A CLOCK KEEP THE CORRECT TIME. WITH WHAT SPEED SHOULD IT MOVE RELATIVE TO OBSERVER SO THAT IT MAY SEEM TO LOSE 4 MIN IN 24 HOURS?
1
Expert's answer
2020-04-10T08:42:23-0400

Solution: The Lorentz transformation for time intervals in different inertial frames of reference is

(1) "\Delta t'=\Delta t\cdot \sqrt{1-\beta^2}" , where "\beta=\frac{V}{c}" , "V" - speed of primed system relative reference frame. Denote "\alpha=\frac{\Delta t'}{\Delta t}" and from (1) find "\beta", we have

(2) "\beta=\sqrt{1-\alpha^2}"

The clock will display the correct time in its own reference system "\Delta t=24hour". In a moving frame of reference it will seem that it is wrong "\Delta t'=24h-4min"

Lets astimate "\alpha=\frac{24h-4min}{24h}=1-\frac{4}{24\cdot 60}=1-2.78\cdot10^{-3}" and denote "\delta=2.78\cdot10^{-3}"

(3) "\alpha^2=(1-\delta)^2=1-2\cdot \delta+\delta^2\simeq1-5.56\cdot 10^{-3}" the square "\delta" can be ignored its value is less than the third significant digit. Substituting (3) in (2) we get

"\beta=\sqrt{1-(1-2\cdot\delta)}=\sqrt{5.56\cdot 10^{-3}}=7.45\cdot 10^{-2}"

The speed of clock should be "V=7.45\cdot 10^{-2}\cdot c=22.4\cdot 10^6 m/s"

Answer: Clock should move relative to observer with speed "2.24\cdot 10^7m/s"